OpenStudy (anonymous):

p+3/m=-1 please help me

OpenStudy (anonymous):

Are you trying to find what p and m are?

OpenStudy (anonymous):

no just what p is

OpenStudy (anonymous):

Ok so you just need to get p by itself then so, \[\frac{ p+3 }{ m }=1\] multiply m to other side p + 3 = m then minus the 3 p = m-3